# SequentialAttention++ for Block Sparsification: Differentiable Pruning Meets Combinatorial Optimization

### v1

@@ 2.1.1 Unnormalized softmax
The softmax operation is a popular differentiable sparsity-inducing technique, where a vector is transformed by exponentiating each entry and normalizing the result. The softmax forms the backbone of many modern ML techniques ranging from multinomial logistic regression to differentiable architecture search Liu et al. (2019) to attention mechanisms and transformers Vaswani et al. (2017), and thus a theoretical understanding of the softmax is critical mission for modern machine learning theory.
We take a step towards this by considering unnormalized softmax, which corresponds to a simple entrywise exponentiation. The unnormalized softmax is a popular alternative to the usual softmax as it still captures its sparsity-inducing properties Amid & Warmuth (2020a, b), while its simplicity allows for more efficient implementations. We show that, in fact, unnormalized softmax can be viewed as a type of log-sum regularization, which is a popular relaxation of the $\lVert\cdot\rVert_{0}$ norm that has been often considered in the machine learning and signal processing literatures Rao & Kreutz-Delgado (1999); Wipf & Nagarajan (2009); Qiao et al. (2020); Tugnait (2022); Zhou et al. (2023).
@@
$\min_{\mathbf{w}\in\mathbb{R}^{t},\boldsymbol{\beta}\in\mathbb{R}^{n}}\mathcal{L}(\{\exp(\mathbf{w}_{i})\boldsymbol{\beta}|_{T_{i}}\}_{i=1}^{t})+\lambda\left\lparen\lVert\mathbf{w}\rVert_{2}^{2}+\lVert\boldsymbol{\beta}\rVert_{2}^{2}\right\rparen$ )
[…]
@@ 2.1.2 $\ell_{1}$-regularized masks
$\min_{\mathbf{w}\in\mathbb{R}^{t},\boldsymbol{\beta}\in\mathbb{R}^{n}}\mathcal{L}(\{\mathbf{w}_{i}\boldsymbol{\beta}|_{T_{i}}\}_{i=1}^{t})+\lambda\left\lparen\lVert\mathbf{w}\rVert_{1}+\lVert\boldsymbol{\beta}\rVert_{2}^{2}\right\rparen$ )
[…]
@@ 2.1.3 Powerpropagation
Finally, we study differentiable pruning techniques that use the network weights themselves as importance scores. The most straightforward implementation of this idea is to square each of the weights, as explored in works such as powerpropagation for neural networks Schwarz et al. (2021), but more complex versions have also been considered Cho et al. (2023). We show how these techniques can be generalized to handle the group setting, and show how they can also be interpreted as an implementation of group sparsity-inducing regularization.
@@
$\min_{\mathbf{w}\in\mathbb{R}^{t},\boldsymbol{\beta}\in\mathbb{R}^{n}}\mathcal{L}(\{\lVert\boldsymbol{\beta}|_{T_{i}}\rVert_{2}\boldsymbol{\beta}|_{T_{i}}\}_{i=1}^{t})+\lambda\lVert\boldsymbol{\beta}\rVert_{2}^{2}$
[…]
@@ 2.2 Unique sparse global minima
$q^{-1}\left\lparen\sum_{i=1}^{t}q(\lVert\boldsymbol{\beta}|_{T_{i}}\rVert_{2})\right\rparen=q^{-1}\left\lparen q(\lVert\boldsymbol{\beta}|_{T_{j}}\rVert_{2})\right\rparen=\lVert\boldsymbol{\beta}|_{T_{j}}\rVert_{2}.$
@@ 3 The SequentialAttention++ algorithm
As discussed, the basic form of SequentialAttention++ is a combination of Sequential Attention and ACDC. More specifically, we first introduce the sequential attention parameterization, where we multiply each selection candidate (e.g. block, in the case of block sparsification) by a trainable attention weight. Then, we employ the alternating compressed and decompressed phases of ACDC on the attention weights (instead of doing so on the selection candidates’ weights themselves). This helps factor out the inherent magnitude differences between model parameters, which do not necessarily correlate with importances. The basic algorithm can be seen in Algorithm 1.
Algorithm 1 Feed-forward layer with the basic version of SequentialAttention++ to select top $k$ parameters from a kernel $\mathbf{W}$. We omit sparsification phases for simplicity.
  function $\text{FF}({{\bf X}\in\mathbb{R}^{b\times n}}:\text{input batch},t:\text{training step})$
     Trainable params:
     Kernel matrix ${\bf W}\in\mathbb{R}^{n\times m}$, Logits ${\bf L}\in\mathbb{R}^{n\times m}$
     ${\bf A}=nm\cdot e^{\bf L}/\sum e^{\bf L}$
     ${\bf\hat{W}}={\bf W}\odot{\bf A}\odot\text{Mask}({\bf A},t)$
     return ${\bf X}{\bf\hat{W}}$
  end function
@@ 3.1 The sparsification phase
One drawback of sparse/dense phases is that the dense-to-sparse transition is abrupt. Since the lowest-magnitude weights are instantly pruned, this neglects correlations between these pruned parameters. If we were to re-train the model after pruning one parameter at a time, the picture could be drastically different, since low-magnitude weights could grow (this could happen e.g. due to parameter redundancy). In fact, this effect was highlighted by Kuznedelev et al. (2023a), who devised a backward selection method based on correlations as captured by the Hessian.
Inspired by this approach, we incorporate a backward selection phase between the dense and sparse phases, which we call the sparsification phase. In this phase, we gradually prune the least important features based on the attention weights. This gradual process allows the model to re-adjust the attention weights after some parameters are pruned. The importance of this phase is validated using ablation experiments in Appendix D.1. We use an exponential pruning schedule, to prune more aggressively in the beginning of the phase, and more carefully at the end (as we approach the desired number of candidates $k$). A comparison of the sparsity schedules of ACDC and SequentialAttention++ can be found in Figure 2. We use the sparsity schedule $\text{sparsity}(t)=s\cdot\frac{1-e^{-ct}}{1-e^{-c}}$ for $t\in[0,1]$, where $s$ is the target sparsity. This interpolates between sparsity 00 and $s$, and constitutes a single sparsification phase. We choose the constant $c=4$ (for an ablation analysis, see Appendix D.2).
Figure 2: Sparsity schedules of ACDC and SequentialAttention++. ACDC uses an instant dense-to-sparse transition, while SequentialAttention++ uses an exponential sparsity schedule.
@@ 4 Experiments
We evaluate our algorithms via sparsification tasks where a dense DNN is approximated by block-sparse counterparts, at various block sizes and sparsities. Our experiments are performed on the ImageNet and Criteo datasets.
ImageNet Deng et al. (2009).
ImageNet is the most widely used vision dataset and is considered as the de facto benchmark in the neural network pruning literature, culminating in the state of the art results in Kuznedelev et al. (2023b). We use ResNet50 and a standard training setup (90 epochs, SGD with cosine learning rate and momentum, weight decay). We reshape the $4$-dimensional ($H\times W\times C_{\text{in}}\times C_{\text{out}}$) kernel tensors used in convolutional layers to 2D matrices of shape $HWC_{\text{in}}\times C_{\text{out}}$, which define the 2D block candidates for pruning. We prune all layers uniformly, except for layers with $<100$ blocks, which we do not prune at all, to avoid degeneracy at high sparsities.
Criteo Diemert et al. (2017).
Criteo is a standard public dataset for the clickthrough rate (CTR) prediction task, which consists of 33M training examples with 13 numerical and 26 categorical features. The model we sparsify is a standard fully connected DNN with three $400$-width layers and an additional embedding layer to transform each input feature into an embedding vector of size $10$ (for a total embedding width of $390$). We prune the first dense layer after the embedding layer. We use Adam optimizer with a learning rate that decays exponentially from $2\cdot 10^{-2}$ to $3\cdot 10^{-4}$. We train to minimize the cross-entropy loss for $25$ epochs with a batch size of $32768$.
We evaluate the performance of our sparsification algorithm at varying sparsities $p$ and block sizes $B$, where a sparsity $p$ indicates that the DNN layer will only have a $1-p$ fraction of nonzero entries, and a block size of $B$ indicates that the nonzero entries are arranged in $B\times B$ blocks. Note that for a fixed sparsity, larger block sizes generally translate to improved efficiency due to improved hardware utilization, but also degrades quality. Block size of $1$ corresponds to unstructured pruning.
@@ 4.1 Baseline algorithms
We compare our SequentialAttention++ algorithm to three other representative prior algorithms for DNN pruning. The first is basic magnitude pruning, which is a popular and effective algorithm where the weights are sparsified by keeping the weights with the largest magnitude after training Frankle & Carbin (2019). We use it in the block setting by keeping the largest blocks in Frobenius norm. The second algorithm is a block generalization of Powerpropagation Schwarz et al. (2021), which combines magnitude pruning with a differentiable pruning technique where sparsity is encouraged by squaring the weights. While the original Powerpropagation algorithm did not handle the block sparsification setting, we show that multiplying each block by the Frobenius norm leads to a provable generalization (see Lemma 2.3). Finally, we consider ACDC Peste et al. (2021), which is an adaptation of iterative hard thresholding (IHT) Blumensath & Davies (2009) to the setting of neural network sparsification, and has produced the state-of-the-art pruning results for ImageNet Kuznedelev et al. (2023b). For all algorithms and datasets, we include a fine-tuning phase at the end of training, using the pruned model, and evaluate the final pruned model on the test set.
@@ 4.2 Results
Appendix A Missing proofs from Section 2
Proof of Lemma 2.1.
Proof of Lemma 2.2.
Proof of Lemma 2.3.
so summing over the groups gives the claimed result. ∎
A.1 Unique sparse global minima
Proof of Theorem 1.1.
[…]
Thus, $\boldsymbol{\beta}^{*}$ must be the unique minimizer of (1). ∎
Appendix B OMPR via nonconvex regularization
We show that our results from Section 2.2 together with recent work of Axiotis & Yasuda (2023) give provable guarantees for a local search algorithm based on orthogonal matching pursuit with replacement using nonconvex regularization.
We first introduce some definitions needed to state our result.
Definition B.1.
Definition B.2 (Restricted strong convexity and smoothness).
We will now obtain provable guarantees for Algorithm 2 in Theorem B.3.
Algorithm 2 OMPR via nonconvex regularization
  Initialize $S$ arbitrarily such that $\lvert S\rvert=k^{\prime}$
  for $i=1,\dots,R$ do
     Let
for $\lambda$ sufficiently large
     $S\leftarrow S\cup\{i\}\setminus\{j\}$
  end for
Theorem B.3 (OMPR via nonconvex regularization).
Proof.
By Theorem 1.1, if the optimization problem in Algorithm 2 with $q$ replaced by the absolute value function has a unique minimizer with group sparsity at most $1$, then $\hat{\boldsymbol{\beta}}$ is a unique global minimizer with group sparsity at most $1$, and coincides with this Group LASSO solution. Lemma 3.2 of Axiotis & Yasuda (2023) then establishes that this solution is supported on the group that maximizes the $\ell_{2}$ norm of the gradient, which in turn implies Theorem B.3 via guarantees for the group orthogonal matching pursuit with replacement algorithm (Corollary A.10 of Axiotis & Yasuda (2023)). ∎
Appendix C Additional details on experiments
C.1 Additional tricks
In addition to the basic algorithm described in Section 3, our implementation of SequentialAttention++ incorporates several other ingredients for improved empirical performance. First, we confirm the observation of Peste et al. (2021) that resetting the optimizer between each phase of SequentialAttention++ is crucial for good performance. We note that this is also suggested by our theoretical results (Theorem B.3), which suggests that each of the dense and sparse phases should be thought of as a separate optimization problem that is solved independently. Similarly to Kuznedelev et al. (2023b), we also observe that weight decay significantly boosts performance, even when applied to the attention logits.

### v2

@@ 2.1.1 Unnormalized softmax
The softmax is a popular differentiable sparsity-inducing technique, where a vector is transformed by exponentiating each entry and normalizing the result. The softmax forms the backbone of many modern ML techniques ranging from multinomial logistic regression to differentiable architecture search (Liu et al., 2019) to attention mechanisms and transformers (Vaswani et al., 2017), and thus a theoretical understanding of the softmax is critical mission for modern machine learning theory.
We take a step towards this by considering unnormalized softmax, which corresponds to a simple entrywise exponentiation. The unnormalized softmax is a popular alternative to the usual softmax as it still captures its sparsity-inducing properties (Amid and Warmuth, 2020a, b), while its simplicity allows for more efficient implementations. We show that, in fact, unnormalized softmax can be viewed as a type of log-sum regularization, which is a popular relaxation of the $\lVert\cdot\rVert_{0}$ norm that has been often considered in the machine learning and signal processing literatures (Rao and Kreutz-Delgado, 1999; Wipf and Nagarajan, 2009; Qiao et al., 2020; Tugnait, 2022; Zhou et al., 2023).
@@
$\min_{\mathbf{w}\in\mathbb{R}^{t},\boldsymbol{\beta}\in\mathbb{R}^{n}}\mathcal{L}(\{\exp(\mathbf{w}_{i})\boldsymbol{\beta}|_{T_{i}}\}_{i=1}^{t})+\lambda\left\lparen\lVert\mathbf{w}\rVert_{2}^{2}+\lVert\boldsymbol{\beta}\rVert_{2}^{2}\right\rparen=\min_{\mathbf{u}\in\mathbb{R}^{n}}\mathcal{L}(\mathbf{u})+\lambda\sum_{i=1}^{t}q(\lVert\mathbf{u}|_{T_{i}}\rVert_{2})$ )
@@ 2.1.2 $\ell_{1}$-regularized masks
$\min_{\mathbf{w}\in\mathbb{R}^{t},\boldsymbol{\beta}\in\mathbb{R}^{n}}\mathcal{L}(\{\mathbf{w}_{i}\boldsymbol{\beta}|_{T_{i}}\}_{i=1}^{t})+\lambda\left\lparen\lVert\mathbf{w}\rVert_{1}+\lVert\boldsymbol{\beta}\rVert_{2}^{2}\right\rparen=\min_{\mathbf{u}\in\mathbb{R}^{n}}\mathcal{L}(\mathbf{u})+\frac{3}{2}2^{1/3}\lambda\sum_{i=1}^{t}\lVert\mathbf{u}|_{T_{i}}\rVert_{2}^{2/3}$
@@ 2.1.3 Powerpropagation
Finally, we study differentiable pruning techniques that use the network weights themselves as importance scores. The most straightforward implementation of this idea is to square each of the weights, as explored in works such as powerpropagation for neural networks (Schwarz et al., 2021), but more complex versions have also been considered (Cho et al., 2023). We show how these techniques can be generalized to handle the group setting, and show how they can also be interpreted as an implementation of group sparsity-inducing regularization.
@@
$\min_{\mathbf{w}\in\mathbb{R}^{t},\boldsymbol{\beta}\in\mathbb{R}^{n}}\mathcal{L}(\{\lVert\boldsymbol{\beta}|_{T_{i}}\rVert_{2}\boldsymbol{\beta}|_{T_{i}}\}_{i=1}^{t})+\lambda\lVert\boldsymbol{\beta}\rVert_{2}^{2}=\min_{\mathbf{u}\in\mathbb{R}^{n}}\mathcal{L}(\mathbf{u})+\lambda\sum_{i=1}^{t}\lVert\mathbf{u}|_{T_{i}}\rVert_{2}$
@@ 2.2 Unique sparse global minima
$q^{-1}\left\lparen\sum_{i=1}^{t}q(\lVert\boldsymbol{\beta}|_{T_{i}}\rVert_{2})\right\rparen=q^{-1}\left\lparen q(\lVert\boldsymbol{\beta}|_{T_{j}}\rVert_{2})\right\rparen=\lVert\boldsymbol{\beta}|_{T_{j}}\rVert_{2}.\qed$
@@ 3 The SequentialAttention++ algorithm
Weight magnitude is a simple and reliable importance score used to prune candidates (in our case, blocks) in a sparse optimization problem. In many cases, however, the magnitudes do not correlate very well with the true importances of the candidates. This has been observed e.g. in Axiotis and Sviridenko (2021, 2022), who showed that the magnitude pruning criterion used in the IHT algorithm is provably suboptimal even for simple sparse regression tasks, and proposed an adaptive weight decay to deal with this issue. One reason for the suboptimality of magnitude pruning is that the weights are not encouraged to be sparse during model training, leading to redundancy. Methods such as Powerpropagation (Schwarz et al., 2021) and Sequential Attention (Yasuda et al., 2023) have been proposed to address this issue by explicitly encoding a non-convexity that encourages weights to be concentrated on a sparse subset (this can be viewed as weight re-parameterization or concave regularization, as shown in Section 2).
To test the hypothesis that softmax attention weights are higher-quality importance scores, we consider one-shot block pruning based on the softmax attention scores used in Sequential Attention (see Figure 1 on how to apply it to blocks), and we compare it with block magnitude (Frobenius norm) pruning. The results in Figure 2(a) suggest that softmax attention scores are generally more reliable as block importance scores, especially for larger block sizes. This leads us to adopt the softmax parameterization in our algorithm.
As observed e.g. in Peste et al. (2021), one-shot pruning approaches are significantly suboptimal compared to iterative pruning approaches such as ACDC. We use a similar alternating compressed and decompressed phases approach as ACDC, but we apply it on the softmax attention weights instead of the block magnitudes. This establishes SequentialAttention++ as a combination between Sequential Attention and ACDC. The basic algorithm can be seen in Algorithm 1.
Algorithm 1 Feed-forward layer with the basic version of SequentialAttention++ to select top $k$ parameters from a kernel $\mathbf{W}$.
@@
  function $\text{FeedForward}({{\bf X}\in\mathbb{R}^{b\times n}}:\text{input batch},t:\text{training step})$
     Trainable params:
     Kernel ${\bf W}\in\mathbb{R}^{n\times m}$, Logits ${\bf L}\in\mathbb{R}^{n\times m}$
     ${\bf A}=nm\cdot e^{\bf L}/\sum e^{\bf L}$
     ${\bf\hat{W}}={\bf W}\odot{\bf A}\odot\text{Mask}({\bf A},t)$
     return ${\bf X}{\bf\hat{W}}$
  end function
@@ 3.1 The sparsification phase
One drawback of sparse/dense (compression/decompression) phases is that the dense-to-sparse transition is abrupt. Since the lowest-magnitude weights are instantly pruned, this neglects correlations between these pruned parameters. If we were to re-train the model after pruning one parameter at a time, the picture could be drastically different, since low-magnitude weights could grow (this could happen e.g. due to parameter redundancy). In fact, this effect was highlighted by Kuznedelev et al. (2023a), who devised a backward selection method based on correlations as captured by the Hessian.
Inspired by this approach, we incorporate a backward selection phase between the dense and sparse phases, which we call the sparsification phase. In this phase, we gradually prune the least important features based on the attention weights. This gradual process allows the model to re-adjust the attention weights after some parameters are pruned. The importance of this phase is validated by ablation experiments in Appendix D.1. We use an exponential pruning schedule, to prune more aggressively in the beginning of the phase, and more carefully at the end (as we approach the desired number of candidates $k$). A comparison of the sparsity schedules of ACDC and SequentialAttention++ can be found in Figure 2(b). We use the sparsity schedule $\text{sparsity}(t)=s\cdot\frac{1-e^{-ct}}{1-e^{-c}}$ for $t\in[0,1]$, where $s$ is the target sparsity. This interpolates between sparsity 00 and $s$, and constitutes a single sparsification phase. We choose the constant $c=4$ (for an ablation analysis, see Appendix D.2).
Validation Accuracy
Block size: $8\times 8$ Block size: $16\times 16$
$70\%$ $80\%$ $90\%$ $95\%$ $70\%$ $80\%$ $90\%$ $95\%$
$-0.12$ $-0.11$ $+0.10$ $+0.19$ $+0.13$ $-0.21$ $-1.33$
Block size: $32\times 32$ Block size: $64\times 64$
$68\%$ $78\%$ $88\%$ $92\%$ $58\%$ $66\%$ $74\%$ $79\%$
$+0.32$ $+0.58$ $+0.71$ $+3.17$ $+2.54$ $+2.81$ $+2.85$ $+5.54$
(a) Quantifying the effectiveness of magnitude vs softmax attention as block importance scores (ResNet50 on ImageNet). For different block sizes and sparsities, we show the difference between the validation accuracy (in percentage points) of a model pruned one-shot based on the softmax attention scores minus one pruned based on the block magnitudes (Frobenius norms). We train dense models for the first half of training, prune once, and then continue to train the remaining blocks for the second half of training. The experimental setup is as in Section 4.
(b) Sparsity schedules of ACDC and SequentialAttention++. ACDC uses an instant dense-to-sparse transition, while SequentialAttention++ uses an exponential sparsity schedule.
Figure 2: (a) Softmax attention vs magnitude pruning, and (b) the Sparsification phase.
@@ 4 Experiments
We evaluate our algorithms on sparsification tasks where a dense DNN is approximated by block-sparse counterparts, at various block sizes $B$ and sparsities $p$, where a sparsity $p$ indicates that the DNN layer will only have a $1-p$ fraction of nonzero entries, and a block size of $B$ indicates that the nonzero entries are arranged in $B\times B$ blocks. Note that for a fixed sparsity, larger block sizes generally translate to improved efficiency due to improved hardware utilization, but also degrades quality. Block size of $1$ corresponds to unstructured pruning. Our experiments are performed on the ImageNet and Criteo datasets. More details on the setup can be found in Section C.1.
@@ 4.1 Baseline algorithms
We compare our SequentialAttention++ algorithm to three other representative prior algorithms for DNN pruning. The first is basic magnitude pruning, which is a popular and effective algorithm where the weights are sparsified by keeping the weights with the largest magnitude after training (Frankle and Carbin, 2019). We use it in the block setting by keeping the largest blocks in Frobenius norm. The second algorithm is a block generalization of Powerpropagation (Schwarz et al., 2021), which combines magnitude pruning with a differentiable pruning technique where sparsity is encouraged by squaring the weights. While the original Powerpropagation algorithm did not handle the block sparsification setting, we show that multiplying each block by the Frobenius norm leads to a provable generalization (see Lemma 2.3). Finally, we consider ACDC (Peste et al., 2021), which is an adaptation of iterative hard thresholding (IHT) (Blumensath and Davies, 2009) to the setting of neural network sparsification, and has produced the state-of-the-art pruning results for ImageNet (Kuznedelev et al., 2023b). For all algorithms and datasets, we include a fine-tuning phase at the end of training, using the pruned model, and evaluate the final pruned model on the test set.
@@ 4.2 Results
5 Conclusion
In this work, we unified, generalized, and improved prior approaches to neural network pruning via a framework which combines differentiable pruning with combinatorial optimization algorithms, in particular local search techniques. Theoretically, we gave a unified analysis of a wide class of existing techniques via a connection to nonconvex regularization, and proved novel properties about sparse convex optimization with nonconvex regularization. In particular, we established natural conditions under which nonconvex regularization yields a unique group-sparse global minimum that is supported on the group that maximizes the $\ell_{2}$ norm of the gradient, thus yielding provable guarantees for group sparse convex optimization. Empirically, we proposed a novel algorithm, SequentialAttention++, which outperforms prior methods on standard benchmark datasets for neural network sparsification.
We conclude with a few open directions which we believe to be interesting for future work. The first is on characterizing the nature of critical points and local minima of nonconvex-regularized convex problems. This would be a more practically useful variation on our result, which only establishes provable guarantees for the global minimizer. For our second question, we ask whether one can theoretically establish that nonconvex regularization yields better optimization guarantees than the LASSO. In our work, we have only shown that the quality of solutions found by nonconvex regularization can match the LASSO for a wide variety of nonconvex regularizers, but we do not theoretically establish that this formulation is better. It would be interesting to show, e.g., that nonconvex regularizers allow for faster convergence to the sparse global minimizer.
Appendix A Missing proofs from Section 2
Proof of Lemma 2.1.
Proof of Lemma 2.2.
Proof of Lemma 2.3.
so summing over the groups gives the claimed result. ∎
A.1 Unique sparse global minima
Proof of Theorem 1.1.
[…]
Thus, $\boldsymbol{\beta}^{*}$ must be the unique minimizer of (1). ∎
Appendix B OMPR via nonconvex regularization
We show that our results from Section 2.2 together with recent work of Axiotis and Yasuda [2023] give provable guarantees for a local search algorithm based on orthogonal matching pursuit with replacement using nonconvex regularization.
We first introduce some definitions needed to state our result.
Definition B.1.
Definition B.2 (Restricted strong convexity and smoothness).
We will now obtain provable guarantees for Algorithm 2 in Theorem B.3.
Algorithm 2 OMPR via nonconvex regularization
  Initialize $S$ arbitrarily such that $\lvert S\rvert=k^{\prime}$
  for $i=1,\dots,R$ do
     Let
for $\lambda$ sufficiently large
     $S\leftarrow S\cup\{i\}\setminus\{j\}$
  end for
Theorem B.3 (OMPR via nonconvex regularization).
Proof.
By Theorem 1.1, if the optimization problem in Algorithm 2 with $q$ replaced by the absolute value function has a unique minimizer with group sparsity at most $1$, then $\hat{\boldsymbol{\beta}}$ is a unique global minimizer with group sparsity at most $1$, and coincides with this Group LASSO solution. Lemma 3.2 of Axiotis and Yasuda [2023] then establishes that this solution is supported on the group that maximizes the $\ell_{2}$ norm of the gradient, which in turn implies Theorem B.3 via guarantees for the group orthogonal matching pursuit with replacement algorithm (Corollary A.10 of Axiotis and Yasuda [2023]). ∎
Appendix C Additional details on experiments
C.1 Experimental setup
ImageNet [Deng et al., 2009].
ImageNet is the most widely used vision dataset and is considered as the de facto benchmark in the neural network pruning literature, culminating in the state of the art results in Kuznedelev et al. [2023b]. We use ResNet50 and a standard training setup (90 epochs, SGD with cosine learning rate and momentum, weight decay). We reshape the $4$-dimensional ($H\times W\times C_{\text{in}}\times C_{\text{out}}$) kernel tensors used in convolutional layers to 2D matrices of shape $HWC_{\text{in}}\times C_{\text{out}}$, which define the 2D block candidates for pruning. We prune all layers uniformly, except for layers with $<100$ blocks, which we do not prune at all, to avoid degeneracy at high sparsities.
Criteo [Diemert et al., 2017].
Criteo is a standard public dataset for the clickthrough rate (CTR) prediction task, which consists of 33M training examples with 13 numerical and 26 categorical features. The model we sparsify is a standard fully connected DNN with three $400$-width layers and an additional embedding layer to transform each input feature into an embedding vector of size $10$ (for a total embedding width of $390$). We note that a simple MLP is often a fairly competitive model for this task [Naumov et al., 2019]. We prune the first dense layer after the embedding layer. We use Adam optimizer with a learning rate that decays exponentially from $2\cdot 10^{-2}$ to $3\cdot 10^{-4}$. We train to minimize the cross-entropy loss for $25$ epochs with a batch size of $32768$.
C.2 Block sparsification results on Criteo
We give our block sparsification results on the Criteo dataset in Table 2.
@@ C.2 Block sparsification results on Criteo
C.3 Additional tricks
In addition to the basic algorithm described in Section 3, our implementation of SequentialAttention++ incorporates several other ingredients for improved empirical performance. First, we confirm the observation of Peste et al. [2021] that resetting the optimizer between each phase of SequentialAttention++ is crucial for good performance. We note that this is also suggested by our theoretical results (Theorem B.3), which suggests that each of the dense and sparse phases should be thought of as a separate optimization problem that is solved independently. Similarly to Kuznedelev et al. [2023b], we also observe that weight decay significantly boosts performance, even when applied to the attention logits.